OpenStudy (anonymous):

Solve for x: "The fourth root of (2x+3) equals 2."

OpenStudy (jdoe0001):

\(\large \bf \sqrt[4]{2x+3}=2\quad \textit{raise both sides to the }^4\)

OpenStudy (mathmale):

^4\sqrt 2x+3 = 2 is not sufficiently clear. You seem to mean "the fourth root of 2x+3) equals 2," so I've taken the liberty of revising your problem statement. Yes: Take the fourth power of each side of the equation. Simplify the result, and solve for x. Be sure to check you answer.